# Improved GUI Grounding via Iterative Narrowing

### v1

@@ Abstract
GUI grounding, the task of identifying a precise location on an interface image from a natural language query, plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework called Iterative Narrowing (IN) to further enhance the performance of both general and fine-tuned models in GUI grounding. For evaluation, we tested our method on a comprehensive benchmark comprising different UI platforms.
@@ 1 Introduction
Instead of training a model, we propose a visual prompting framework called Iterative Narrowing (IN) to improve the performance of current ones. Our approach treats the model’s initial position predictions as an approximation. The framework iteratively refines the prediction by focusing on progressively smaller, cropped regions that center around the previous prediction. This refinement process can be repeated for multiple iterations.
Figure 1: A visual demonstration of IN. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
@@ 4.1 Evaluation Results
Table 1: Overall average accuracy comparing baseline against our method (IN) on the ScreenSpot benchmark.
@@
Baseline IN Baseline IN
@@
Baseline IN Baseline IN
@@
Baseline IN Baseline IN
@@ 5 Conclusion
In this work, we introduced Iterative Narrowing (IN), a visual prompting framework designed to enhance the GUI grounding capabilities of Vision-Language Models (VLMs). Through iterative refinement of predictions, IN improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface. Evaluation on the ScreenSpot [4] benchmark demonstrated that IN improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [7], where substantial improvements were observed compared to baselines. However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.

### v2

@@ Abstract
Graphical User Interface (GUI) grounding plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework that employs an iterative narrowing mechanism to improve the performance of both general and fine-tuned models in GUI grounding by up to 61%. For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@ 1 Introduction
Instead of training a model, we propose a visual prompting framework that employs an iterative narrowing mechanism to improve the grounding performance of current VLMs. Our approach treats the model’s initial position predictions as an approximation. The framework iteratively refines the prediction by focusing on progressively smaller, cropped regions that center around the previous prediction. This refinement process can be repeated for multiple iterations.
Figure 1: A visual demonstration of our method. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
@@ 2.2 GUI Agents
2.3 Iterative Visual Reasoning Beyond Convolutions
2.4 Iterative Localization in Proposal-Free Referring Expression Grounding
Sun et al. [7] tackled proposal-free referring expression grounding introducing a similar mechanism called iterative shrinking, which is guided by a reinforcement learning agent to localize target objects based on query sentences.
This iterative approach parallels our method, though Sun et al. uses reinforcement learning, while our method relies on VLM predictions to refine the target.
@@ 4.1 Evaluation Results
Table 1: Overall average accuracy comparing baseline against our method on the ScreenSpot benchmark.
@@
Baseline Ours Baseline Ours
@@
Baseline Ours Baseline Ours
@@
Baseline Ours Baseline Ours
@@ 5 Conclusion
In this work, we introduced a visual prompting framework designed to perform an iterative narrowing mechanism. As a result, this enhances the GUI grounding capabilities of Vision-Language Models (VLMs). Through iterative refinement of predictions, our method improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface. Evaluation on the ScreenSpot [4] benchmark demonstrated that our method improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [8], where substantial improvements were observed compared to baselines. However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.